OpenStudy (anonymous):

When graphed, the circle with equation x^2 +y^2 +4x -4y +7 =0 will lie ENTIRELY in Quadrant A) I. B) II. C) III. D) IV.

jimthompson5910 (jim_thompson5910):

you can use desmos to graph https://www.desmos.com/calculator

OpenStudy (anonymous):

I have a standardized test coming up on Tuesday and I will not be able to use the website you recommended. Could you tell me the steps in graphing this equation?

OpenStudy (triciaal):

alternate group the x portion group the y portion find the value needed to complete each square the center of the circle is (h,k) using (x-h)^2 + (y-k)^2 = r^2 r is the radius of the circle

jimthompson5910 (jim_thompson5910):

x^2 +y^2 +4x -4y +7 =0 (x^2+4x) +(y^2 -4y) +7 =0 ------------------------------------------------------- focus on just the (x^2+4x) terms for now. What is needed here to complete the square?

OpenStudy (anonymous):

a 4 is needed?

jimthompson5910 (jim_thompson5910):

correct, you take half of the x coefficient 4 to get 2 then you square 2 to get 4 we add and subtract 4 so things stay balanced (effectively it's like adding 0) x^2 + 4x x^2 + 4x + 4 - 4 (x^2 + 4x + 4) - 4 (x+2)^2 - 4

jimthompson5910 (jim_thompson5910):

so x^2 + 4x turns into (x+2)^2 - 4

jimthompson5910 (jim_thompson5910):

what does y^2 - 4y turn into when we complete the square?

OpenStudy (anonymous):

(y-2)^2 + 4

jimthompson5910 (jim_thompson5910):

let's expand (y-2)^2 + 4 out (y-2)^2 + 4 (y-2)(y-2) + 4 y(y-2) - 2(y-2) + 4 y^2 - 2y - 2y + 4 + 4 y^2 - 4y + 8 so we went from (y-2)^2 + 4 to y^2 - 4y + 8 but we wanted to go from (y-2)^2 + 4 to y^2 - 4y

jimthompson5910 (jim_thompson5910):

y^2 - 4y y^2 - 4y + 4 - 4 (y^2 - 4y + 4) - 4 (y-2)^2 - 4 If you expanded (y-2)^2 - 4 out and simplified, you would get y^2 - 4y again

OpenStudy (anonymous):

So sorry! I've mistaken my -4 as +4

jimthompson5910 (jim_thompson5910):

that's ok

jimthompson5910 (jim_thompson5910):

after those completing the square steps for x and y, we will have x^2 +y^2 +4x -4y +7 =0 (x^2+4x) +(y^2 -4y) +7 =0 (x+2)^2 - 4 + (y-2)^2 - 4 = 0 now you must get the equation into (x-h)^2 + (y-k)^2 = r^2 form

OpenStudy (anonymous):

wait, where did the 7 go?

jimthompson5910 (jim_thompson5910):

oh my bad lol

jimthompson5910 (jim_thompson5910):

x^2 +y^2 +4x -4y +7 =0 (x^2+4x) +(y^2 -4y) +7 =0 (x+2)^2 - 4 + (y-2)^2 - 4 +7 = 0

OpenStudy (anonymous):

would I write it like this: (x+2)^2 + (y-2)^2= 1 ?

jimthompson5910 (jim_thompson5910):

that is correct

jimthompson5910 (jim_thompson5910):

what is the center of that circle?

OpenStudy (anonymous):

Thanks so much I know how to graph it from here!

jimthompson5910 (jim_thompson5910):

you're welcome